what is the perimeter of a triangle which sides measure 6,8,and 10 inches?
Answer by jim_thompson5910(35256)

Perimeter of a triangle:

P=side1+side2+side3

P=6+8+10=14+10=24


So the perimeter of the triangle is 24 inches